# Zephyrus: Scaling Gateways Beyond the Petabit-Era with DPU-Augmented Hierarchical Co-Offloading

### v1

@@ 4.2 Packet Processing across Heterogeneous Hardware
4.3 Mechanisms for Consistent and Efficient Heterogeneous P4 Pipeline
@@ 5.1 Performance Comparison
Figure 7: Performance comparison of three cloud gateway versions
@@
Figure 8: Latency distribution with different packet numbers
@@
Figure 9: PHV usage across pipelines (the worst case)
@@
Figure 10: Total on-chip resource usage across pipelines
@@
Figure 11: SRAM usage over stages across pipelines
Figure 12: SRAM saturation percentage across pipelines
Figure 13: Table capacity of different versions
5.2 Resource Usage
@@
6 Related Work
In recent years, Data Processing Units (DPUs) have rapidly emerged as a key enabler for cloud and datacenter infrastructure. Recent work highlights their diverse capabilities, including storage offloading [52], near-storage secure filtering [53], and 6G user-plane acceleration [54], demonstrating their growing role in different fields. Sirius [55] further demonstrates this trend by disaggregating a broad class of stateful network functions into pools of P4-programmable NIC cards.
In the context of cloud gateways, previous state-of-the-art works primarily focus on using FPGAs [4, 3] and/or switching ASICs [13, 4] for acceleration. Sailfish [13] uses Tofino ASICs to accelerate forwarding, employing a series of hardware and software co-designs to address the on-chip memory limitations of the Tofino chip. LuoShen [4] leverages FPGAs as an intermediate layer between CPUs and Tofinos to handle complex high-bandwidth forwarding, achieving 1.2 Tbps throughput. Albatross [3], in contrast, is based on FPGAs and x86 CPUs, delivering FPGA-based packet-level load balancing to the host without a switching ASIC.
In contrast, as a centralized gateway, Zephyrus adopts DPUs to augment the Tofino pipeline. This design significantly expands effective table capacity, overcoming the resource shortage limitations in architectures such as Sailfish and Luoshen. Meanwhile, Zephyrus’s three-stage "sandwich" architecture improves its throughput to 1.6 Tbps, a 33% improvement over Luoshen and a 100% improvement over Albatross. To overcome memory limitations, Albatross, which is based on FPGA, suffers a significant performance regression in throughput (4x), packet rate (18x) and latency (10x) compared to Tofino-based gateways. By comparison, Zephyrus uses a DPU-augmented architecture to overcome resource limitations; by hierarchically co-offloading over 99% of packets to hardware, it achieves only a slight performance loss compared to Tofino-based gateways.
As cloud workloads diversify and traffic scales non-linearly, a growing number of hyperscalers are exploring heterogeneity in the network dataplane [56, 57]. However, this very heterogeneity introduces significant development and operational complexity. Programming these disparate hardware targets requires specialized teams and fragmented toolchains, leading to long and costly development cycles. We introduce a unified P4 pipeline to accelerate development cycles [58], significantly simplifying the development and operational complexity.
7 Experiences
7.1 Development and Deployment Experiences
@@ 7.2 Lessons Learned
Conduct thorough evaluation before development. We first evaluated our growth in production traffic and recognized the need to redesign the next generation gateway. Through a systematic evaluation of hardware options, we selected DPUs over FPGAs. While both provide hardware acceleration, DPUs offer significantly lower development complexity. Prior work shows that FPGA-based data plane designs often require tens of thousands of lines of Verilog or HLS code for packet parsing and control logic (e.g., FpgaNIC [59], NetFPGA [60]), while DPU development leverages C/C++ with DPDK libraries or P4 for specific modules, typically implemented in only a few thousand lines of code [61, 62]. Guided by workload analysis including traffic composition, frequency distributions, and resource demands, we adopted a hierarchical forwarding architecture. We also evaluated the expertise within our team and assigned responsibilities accordingly before starting implementation. This careful upfront evaluation directly contributed to reduced TCO, and substantially fewer man-months compared to a less structured development process.
Make the system future-proof. Fueled by continuous breakthroughs in SerDes (Serializer/Deserializer) [63] technology, the aggregate throughput of switching ASICs has followed an exponential growth trajectory, doubling approximately every two years. We implemented and deployed Zephyrus using Intel Tofino switches, which were the state-of-the-art P4-programmable ASICs when our project began. As a long-running system, we must anticipate a future where our existing inventory of a specific ASIC model becomes depleted during large-scale deployment, and the broader hardware landscape will have evolved. Designing an abstract architecture decoupled from specific hardware, building on hardware types that constantly advance, and maintaining an experienced developing team enable us to rapidly migrate to more performant or cost-effective switching ASICs (e.g., Xsight Labs Unveils X1 [64], Broadcom Tomahawk [63]) with minor redevelopment. Meanwhile, in terms of DPUs, the Intel roadmap projects that its upcoming IPUs (a DPU equivalent design) will reach 800 Gbps in 2025 [65], offering strong evidence in favor of our design and showing that Zephyrus is well positioned for future evolution.
7.3 Broader Design Principles and Insights
@@
The principle of software-hardware asymmetry. Our experiences with Zephyrus suggest that architectures can benefit from an intentional software-hardware asymmetry [66]. This experience indicates that hardware’s role is often best suited for low-latency, high-throughput, and repetitive data-plane tasks (e.g., packet parsing, encryption, and basic forwarding). In contrast, software appears well suited for managing complexity and evolving business logic (e.g., policy orchestration, cross-domain optimization, and anomaly detection [67, 68, 69]). Zephyrus’s architecture is a direct embodiment of this principle, combining the Tbps-scale throughput of its ASIC data plane with the rich flexibility of its software components.
An explicit line between standardization and customization. Running a system like Zephyrus at scale forces us to confront a classic tension: how do we keep the system stable while constantly adding new features for different customers? Our answer was to draw an explicit line between standardization and customization. In our observation, core infrastructure components, such as routing protocols, hardware interfaces, and baseline control logic, benefit from conforming to industry standards to promote interoperability [70] and stability. Programmability, in turn, is leveraged at the service edge to enable fine-grained, tenant-specific features (e.g., differentiated QoS policies, dynamic filtering rules). This separation allows us to satisfy unique requirements while minimizing impact on the standardized core, thereby preserving system-wide consistency and robustness.
8 Conclusion

### v2

@@ 4.2. Packet Processing across Heterogeneous Hardware
4.3. Mechanisms for Consistent and Efficient Heterogeneous P4 Pipeline
@@ 5.1. Performance Comparison
Figure 7. Performance comparison of three cloud gateway versions
@@
Figure 8. Latency distribution with different packet numbers
@@
Figure 9. PHV usage across pipelines (the worst case)
@@
Figure 10. Total on-chip resource usage across pipelines
@@
Figure 11. SRAM usage over stages across pipelines
Figure 12. SRAM saturation percentage across pipelines
Figure 13. Table capacity of different versions
5.2. Resource Usage
@@
6. Related Work
In recent years, Data Processing Units (DPUs) have rapidly emerged as a key enabler for cloud and datacenter infrastructure. Recent work highlights their diverse capabilities, including storage offloading (zhao2024tianmen), near-storage secure filtering (batsoyol2025dpuf), and 6G user-plane acceleration (lee2024upf), demonstrating their growing role in different fields. Sirius (bansal2023disaggregating) further demonstrates this trend by disaggregating a broad class of stateful network functions into pools of P4-programmable NIC cards.
In the context of cloud gateways, previous state-of-the-art works primarily focus on using FPGAs (pan2024luoshen; lu2025albatross) and/or switching ASICs (pan2021sailfish; pan2024luoshen) for acceleration. Sailfish (pan2021sailfish) uses Tofino ASICs to accelerate forwarding, employing a series of hardware and software co-designs to address the on-chip memory limitations of the Tofino chip. LuoShen (pan2024luoshen) leverages FPGAs as an intermediate layer between CPUs and Tofinos to handle complex high-bandwidth forwarding, achieving 1.2 Tbps throughput. Albatross (lu2025albatross), in contrast, is based on FPGAs and x86 CPUs, delivering FPGA-based packet-level load balancing to the host without a switching ASIC.
In contrast, as a centralized gateway, Zephyrus adopts DPUs to augment the Tofino pipeline. This design significantly expands effective table capacity, overcoming the resource shortage limitations in architectures such as Sailfish and Luoshen. Meanwhile, Zephyrus’s three-stage ”sandwich” architecture improves its throughput to 1.6 Tbps, a 33% improvement over Luoshen and a 100% improvement over Albatross. To overcome memory limitations, Albatross, which is based on FPGA, suffers a significant performance regression in throughput (4x), packet rate (18x) and latency (10x) compared to Tofino-based gateways. By comparison, Zephyrus uses a DPU-augmented architecture to overcome resource limitations; by hierarchically co-offloading over 99% of packets to hardware, it achieves only a slight performance loss compared to Tofino-based gateways.
As cloud workloads diversify and traffic scales non-linearly, a growing number of hyperscalers are exploring heterogeneity in the network dataplane (Catapult2018). However, this very heterogeneity introduces significant development and operational complexity. Programming these disparate hardware targets requires specialized teams and fragmented toolchains, leading to long and costly development cycles. We introduce a unified P4 pipeline to accelerate development cycles (Bosshart:P4:2014), significantly simplifying the development and operational complexity.
7. Experiences
7.1. Development and Deployment Experiences
@@ 7.2. Lessons Learned
Conduct thorough evaluation before development. We first evaluated our growth in production traffic and recognized the need to redesign the next generation gateway. Through a systematic evaluation of hardware options, we selected DPUs over FPGAs. While both provide hardware acceleration, DPUs offer significantly lower development complexity. Prior work shows that FPGA-based data plane designs often require tens of thousands of lines of Verilog or HLS code for packet parsing and control logic (e.g., FpgaNIC (fpganic_atc22), NetFPGA (netfpga_p4)), while DPU development leverages C/C++ with DPDK libraries or P4 for specific modules, typically implemented in only a few thousand lines of code (dpdk_doca; dpdk_on_dpu). Guided by workload analysis including traffic composition, frequency distributions, and resource demands, we adopted a hierarchical forwarding architecture. We also evaluated the expertise within our team and assigned responsibilities accordingly before starting implementation. This careful upfront evaluation directly contributed to reduced TCO, and substantially fewer man-months compared to a less structured development process.
Make the system future-proof. Fueled by continuous breakthroughs in SerDes (Serializer/Deserializer) technology, the aggregate throughput of switching ASICs has followed an exponential growth trajectory, doubling approximately every two years. We implemented and deployed Zephyrus using Intel Tofino switches, which were the state-of-the-art P4-programmable ASICs when our project began. As a long-running system, we must anticipate a future where our existing inventory of a specific ASIC model becomes depleted during large-scale deployment, and the broader hardware landscape will have evolved. Designing an abstract architecture decoupled from specific hardware, building on hardware types that constantly advance, and maintaining an experienced developing team enable us to rapidly migrate to more performant or cost-effective switching ASICs (e.g., Xsight Labs Unveils X1, Broadcom Tomahawk) with minor redevelopment. Meanwhile, in terms of DPUs, the Intel roadmap projects that its upcoming IPUs (a DPU equivalent design) will reach 800 Gbps in 2025 (kennedy2022ipu), offering strong evidence in favor of our design and showing that Zephyrus is well positioned for future evolution.
7.3. Broader Design Principles and Insights
@@
The principle of software-hardware asymmetry. Our experiences with Zephyrus suggest that architectures can benefit from an intentional software-hardware asymmetry. This experience indicates that hardware’s role is often best suited for low-latency, high-throughput, and repetitive data-plane tasks (e.g., packet parsing, encryption, and basic forwarding). In contrast, software appears well suited for managing complexity and evolving business logic (e.g., policy orchestration, cross-domain optimization, and anomaly detection (openflow08; b4_13; sonata18)). Zephyrus’s architecture is a direct embodiment of this principle, combining the Tbps-scale throughput of its ASIC data plane with the rich flexibility of its software components.
An explicit line between standardization and customization. Running a system like Zephyrus at scale forces us to confront a classic tension: how do we keep the system stable while constantly adding new features for different customers? Our answer was to draw an explicit line between standardization and customization. In our observation, core infrastructure components, such as routing protocols, hardware interfaces, and baseline control logic, benefit from conforming to industry standards to promote interoperability (Saltzer1984) and stability. Programmability, in turn, is leveraged at the service edge to enable fine-grained, tenant-specific features (e.g., differentiated QoS policies, dynamic filtering rules). This separation allows us to satisfy unique requirements while minimizing impact on the standardized core, thereby preserving system-wide consistency and robustness.
8. Conclusion